Determine whether each triangle with sides of given lengths is a right angle
Problem=
21 ft, 28 ft, 35, ft

Answer by mananth(16946)

If 21^2+28^2=35^2 then it is a right triangle
441+784=1225
35^2=1225
By pythagoras theorem this is a right triangle